OpenStudy (anonymous):

x + 5x + 3x + 30 = 180

OpenStudy (angela210793):

9x=150 x=150/9=50/3

OpenStudy (anonymous):

is that rite

OpenStudy (angela210793):

yes it is

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

18

hero (hero):

What happened to your other questions Mido?

OpenStudy (anonymous):

n one answered them:(

hero (hero):

I was going to help you with them